# Explaining Tournament Solutions with Minimal Supports

### v2

@@ 4.1 Top Cycle and Uncovered Set
Let $\mathcal{X}=(C,E_{\mathcal{X}})$ be a MS for $w\in\operatorname{TC}(G)$. Let us now show that every candidate different from $w$ has an in-degree of 1 in $\mathcal{X}$ and that $w$ has an in-degree of 0. According to Proposition 1, for all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a directed path $p_{c}$ from $w$ to $c$ in $\mathcal{X}$. Suppose that exists $c^{*}\in\mathcal{C}\setminus{\{w\}}$ such that there exist distinct $c^{\prime},c^{\prime\prime}$ such that $\{(c^{\prime},c^{*}),(c^{\prime\prime},c^{*})\}\subseteq E_{\mathcal{X}}$. Then $p_{c^{\prime}}+(c^{\prime},c^{*})$ and $p_{c^{\prime\prime}}+(c^{\prime\prime},c^{*})$ are two distinct directed paths from $w$ to $c$ in $\mathcal{X}$. Contradiction. Hence, for all $c\in\mathcal{C}\setminus{\{w\}}$, every candidate different from $w$ has an in-degree of 1 in $\mathcal{X}$. Suppose that exists $c^{*}\in\mathcal{C}\setminus{\{w\}}$ such that $(c^{*},w)\in E_{\mathcal{X}}$. No $p_{c}$ contains $(c^{*},w)$ otherwise as they start at $w$, they would pass by $w$ twice. Let $\mathcal{Y}=(\mathcal{C},E_{\mathcal{Y}})$ be such that $E_{\mathcal{Y}}=E_{\mathcal{X}}\setminus{\{(c^{*},w)\}}$. For all $c\in\mathcal{C}\setminus{\{w\}}$, $p_{c}$ is a directed path from $w$ to $c$ in $\mathcal{Y}$ Thus $w\in\text{NW}_{\operatorname{TC}}(\mathcal{Y})$ but $\mathcal{Y}\subsetneq\mathcal{X}$ so $\mathcal{X}$ is not a MS for $w\in\operatorname{TC}(G)$. Contradiction. Hence, $w$ as an in-degree of 0 in $\mathcal{X}$. Finally, since the sum of the in-degree of all nodes is the number of edges in a graph, $|\mathcal{X}|=m-1$. ∎
Finally, computing a shortest-paths SMS for $\operatorname{TC}$, can be achieved in polynmial time with Dijkstra’s algorithm.
Proposition 3.
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{TC}(G)$, there exists an algorithm which computes a shortest-paths SMS for $w\in\operatorname{TC}(G)$ in $\mathcal{O}(m^{2})$.
Like $\operatorname{TC}$, the uncovered set ($\operatorname{UC}$) has a characterization based on the existence of directed paths between the winner and the other candidates, except that these paths have a maximum length of two. Hence, the following results are derived from the previous propositions.
Corollary 4.
Given a complete tournament $G=(\mathcal{C},E)$, and a winning candidate $w\in\operatorname{UC}(G)$, for all MS $\mathcal{X}$ for $w\in\operatorname{UC}(G)$, $\mathcal{X}$ is a $w$-rooted out-tree of depth at most 2, i.e., for all $c\in\mathcal{C}\setminus{\{w\}}$, there exists a unique directed path of length at most 2 from $w$ to $c$ in $\mathcal{X}$.
Corollary 5.
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{UC}(G)$, for all SMSs $\mathcal{X}$ for $w\in\operatorname{UC}(G)$, we have $|\mathcal{X}|{=}m-1$.
Corollary 6.
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{UC}(G)$, there exists an algorithm which computes a shortest-paths SMS for $w\in\operatorname{UC}(G)$ in $\mathcal{O}(m^{2})$.
4.2 Weighted Uncovered Set
To construct SMSs for the weighted uncovered set ($\operatorname{wUC}$), we generalize the alternative characterization of the uncovered set to the case of weighted tournaments (Döring and Peters 2023, Lemma 2.2).
Lemma 7.
From this lemma, we obtain a similar result to Proposition 4 on the structure of any minimal support for the weighted uncovered set.
Proposition 8.
Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$, a winning candidate $w\in\operatorname{wUC}(G)$, for all MS $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ for $w\in\operatorname{wUC}(G)$, $\mathcal{X}$ is a $w$-rooted out-tree of depth at most 2 such that for all losing candidates $c\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(w,c)\geq\lceil(n+1)/2\rceil$ or exists a unique $c^{\prime}$ such that $\mu_{\mathcal{X}}(w,c^{\prime})+\mu_{\mathcal{X}}(c^{\prime},c)\geq n+1$.
Even if MSs for $\operatorname{wUC}$ follow a similar structure to the unweighted case, we now show that computing a smallest MS is NP-complete.
Proposition 9.
Given $n\geq 2$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$, a winning candidate $w\in\operatorname{wUC}(G)$, and an integer $k$, determining if there exists a SMS of size $k$ is NP-complete.
Proof sketch.
To prove NP-hardness, we proceed with a reduction from the NP-complete set cover problem (Karp 1972). Given an instance of set cover problem with a universe $U$ of $p$ elements and a set $S$ of $q$ subsets of $U$, we show that there exists a set cover of size less than $k\in\mathbb{N}$ if and only if there exists a SMS for $\operatorname{wUC}$ of size less than $k^{\prime}=p+q+k$ for a tournament build from $U$ and $S$. ∎
Even if computing SMSs for $\operatorname{wUC}$ is hard, we provide bounds on their size.
Proposition 10.
Given $n\geq 2$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m\geq 3$, and a winning candidate $w\in\operatorname{wUC}(G)$, for all SMSs $\mathcal{X}$ for $w\in\operatorname{wUC}(G)$, we have $n+m-2{\leq}|\mathcal{X}|{\leq}(n-1)(m-1)$.
4.3 Maximin
Maximin ($\operatorname{MM}$) select the set of candidate with the best worst performance against another alternative. Thus, an MS for $\operatorname{MM}$ must guarantee minimum performance for the winner while ensuring that the other candidates perform less well.
Lemma 11.
Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$, and a winning candidate $w\in\operatorname{MM}(G)$ with a maximin score of $\sigma_{w}$, for all MSs $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ for $w\in\operatorname{MM}(G)$, exists $t\leq\left\lceil n/2\right\rceil$ such that, for all $c\in\mathcal{C}\setminus{\{w\}}$, $\mu_{\mathcal{X}}(w,c)\geq t$ and exists $c^{\prime}\in\mathcal{C}$ such that $\mu_{\mathcal{X}}(c^{\prime},c)\geq n-t$.
From the previous characterization, we can derive a closed-form expression for the size of a SMS for $\operatorname{MM}$.
Proposition 12.
Given $n$ voters, a complete $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{MM}(G)$ with a maximin score of $\sigma_{w}$, for all SMSs $\mathcal{X}$ for $w\in\operatorname{MM}(G)$, let $t=\min(\sigma_{w},\left\lfloor n/2\right\rfloor)$ we have $|\mathcal{X}|{=}n(m-1)-t|\{c:c\in\mathcal{C},\,\mu(w,c)\geq n-t\}|$.
Proof.
@@ 4.3 Maximin
According to Lemma 11, for each candidate different from $w$, n=t+ntn=t+n-t pairwise comparisons are needed. However, one can optimize the size of the MS when $w$ wins ntn-t pairwise comparisons against an opponent $c$ by covering both constraints at once. Indeed, let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ be a SMS for $w\in\operatorname{MM}(G)$ and $t\leq\lfloor n/2\rfloor$. For all $c\in\mathcal{C}$ such that $\mu(w,c)\geq n-t$, if we have $\mu_{\mathcal{X}}(w,c)=n-t$, then for all completions $\mathcal{X}^{\prime}$ of $\mathcal{X}$, we have $\mu_{\mathcal{X}^{\prime}}(w,c)\geq n-t\geq t$. For all the other candidates $c\in\mathcal{C}$ such that $\mu(w,c)<n-t$, we take $\mu_{\mathcal{X}}(w,c)=t$ and $\mu_{\mathcal{X}}(c^{\prime},c)=n-t$ for some candidate $c^{\prime}\in\mathcal{C}$ such that $\mu(c^{\prime},c)\geq n-t$. Then for all completions $\mathcal{X}^{\prime}$ of $\mathcal{X}$, we have $\mu_{\mathcal{X}^{\prime}}(w,c)\geq t$ and $\mu_{\mathcal{X}^{\prime}}(c^{\prime},c)\geq n-t$. Let $t=\min(\sigma_{w},\left\lfloor n/2\right\rfloor)$, for all losing candidate $c\in\mathcal{C}\setminus{\{w\}}$, either $\mu(w,c)\geq n-t$ and only ntn-t pairwise comparisons are needed else n=nt+tn=n-t+t comparisons are needed. With $k=|\{c:c\in\mathcal{C},\,\mu(w,c)\geq n-t\}|$, we have $|\mathcal{X}|=(n-t)k+n(m-1-k)=n(m-1)-tk$. ∎
We then provide bounds on the size of a SMS.
@@ 4.4 Borda and Copeland Rules
The Borda rule ($\operatorname{BO}$) chooses the set of candidates with a maximal total amount of victories across all pairwise comparisons with the other candidates.
We show that the size of a SMS for $\operatorname{BO}$ only depends on the Borda score of the winner and its worst performance against the other candidates.
Proposition 15.
@@
Proof sketch.
We show that there exist three constraints on SMSs inducing three lower bounds on their size, one for the first case and two for the second. Additionally, we prove that each lower bound is binding in its specific sub-case. ∎
From the characterization of the size of SMS for $\operatorname{BO}$, we derive upper and lower bounds on SMSs size.
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$. For all SMSs $\mathcal{X}$ for $w\in\operatorname{BO}(G)$ $\left\lceil\frac{n(m-1)^{2}}{m}\right\rceil\leq|\mathcal{X}|\leq(m-1)\left\lfloor n\frac{m-1}{2}\right\rfloor$.
@@
Finally, we show that all SMSs share a similar structure centered around the winner’s victories by bounding the difference in terms of win count between SMSs.
@@
Since the Copeland rule coincides with the Borda rule on 1-weighted tournaments. The following results are a special case of the previous results obtained on Borda.
@@
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$ and a winning candidate $w\in\operatorname{CO}(G)$ with a Copeland score of $\sigma_{w}$, for all SMS $\mathcal{X}$ for $w\in\operatorname{CO}(G)$:
@@
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$ and a winning candidate $w\in\operatorname{CO}(G)$, for all SMS $\mathcal{X}$ for $w\in\operatorname{CO}(G)$ $m-1\leq|\mathcal{X}|\leq(m-1)\left\lfloor\frac{m-1}{2}\right\rfloor$.
@@ 5 Certified Explanations
  • $a$ is not covered by $b$ since $a$ is preferred to $b$ ($(a,b)\in\mathcal{X}$)
  • $a$ is not covered by $c$ since $a$ is preferred to $b$ ($(a,b)\in\mathcal{X}$) and $b$ is preferred to $c$ ($(b,c)\in\mathcal{X}$)
  • $a$ is not covered by $d$ since $a$ is preferred to $d$ ($(a,d)\in\mathcal{X}$)
@@
  • $a$ wins at least 2 pairwise comparisons in each head-to-head ($\mu_{\mathcal{X}}(a,b)=3$, $\mu_{\mathcal{X}}(a,c)=2$, $\mu_{\mathcal{X}}(a,d)=3$)
  • $b$ wins at most 2=5-3 pairwise comparisons against $a$ ($\mu_{\mathcal{X}}(a,b)=3$)
  • $c$ wins at most 2=5-3 pairwise comparisons against $b$ ($\mu_{\mathcal{X}}(b,c)=3$)
  • $d$ wins at most 2=5-3 pairwise comparisons against $a$ ($\mu_{\mathcal{X}}(a,d)=3$)
Figure 4: From a tournament (a) (resp. (e)), we compute a minimal support for $a\in\operatorname{UC}(G)$ (b) (resp. $a\in\operatorname{MM}(G)$ (f)), we identify its underlying structure (c) (resp. (g)) and we produce a textual explanation (d) (resp. (h)). Examples for the remaining tournament solutions are available in the technical appendix.
We now inspect MSs in light of the four Gricean principles underpinning effective explanations that we sketched in the introduction: quantity, quality, relation and manner (Grice 1975). The formal roots of MSs ensure that the reasoning is sound, thus covering quality. By definition, MSs provide the minimal information necessary to ensure a designated candidate is among the election winners. Thus, MSs are informative enough to support our goal and verify both the quantity and relation principles. Regarding manner, or the way information is delivered, we observe first that our efficient algorithms to compute the smallest MSs for five out of the six studied solutions, ensure that a relatively small amount of information is necessary. However, MSs still suffer from two flaws that limit their explanatory power. First, MSs are (multi-)sets of pairwise comparisons, with a flat structure that makes it hard to gain insights from the unfolding of the tournament solution. Second, their formal nature renders them more suited to be handled by a machine or by domain experts than end-users. We propose to address this by capitalizing on our structural results for MSs to organize the explanation, and by showcasing the production of intuitive textual explanation from MSs.
Figure 4 showcases two examples of this process. For path-based rules such as $\operatorname{UC}$, MSs can be presented as a rooted out-trees encompassing all directed paths from the winner to the other candidates (Figure 4(c)), with each path corresponding to a bullet point in the textual explanation (see Figure 4(d)). For myopic-score rules such as $\operatorname{MM}$, MSs can present the neighbourhoods of each candidate (see Figure 4(g)) from which candidates’ score can be bounded to create a compact textual explanation (see Figure 4(h)).
@@ 6 Conclusions
Minimal supports are compact and can serve to provide certified explanations for tournament solutions, but they also have far-reaching connections with several classical problems in computational social choice. In contrast with previous results on the margin of victory (Brill, Schmidt-Kraepelin, and Suksompong 2020; Döring and Peters 2023) for the top cycle, the uncovered set, Copeland, and Borda, we prove that the size of a smallest minimal support is the same for all candidates in the winner set, preventing its use as a refinement of those tournament solutions. From a query complexity perspective, while the top cycle and the uncovered set require $\Omega(m^{2})$ queries (Maiti and Dey 2024, Theorem 3.5 & 3.8), we show that, with a perfect oracle, exactly m1m-1 queries are needed. Although finding the smallest minimal explanation is generally hard, we provide algorithms for the top cycle, the uncovered set, Copeland, maximin and Borda with complexity $\mathcal{O}(m^{2}\log n)$, which corresponds to the size of a tournament.
Looking ahead, several key directions merit further exploration: empirically evaluating the usefulness of the presented explanations, studying a broader range of rules, and analyzing structural properties of minimal supports. Finally, in an approach similar to that of De Donder, Le Breton, and Truchon (2000), minimal supports for different rules on the same tournament could be compared from a set-theoretical point of view.
7 Acknowledgment
Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work is supported by ERC grant 101166894 “Advancing Digital Democratic Innovation” (ADDI).

### v3

@@ 4.3 Maximin
According to Lemma 11, for each candidate different from $w$, n=t+ntn=t+n-t pairwise comparisons are needed. However, one can optimize the size of the MS when $w$ wins ntn-t pairwise comparisons against an opponent $c$ by covering both constraints at once. Indeed, let $\mathcal{X}=(\mathcal{C},\mu_{\mathcal{X}})$ be an SMS for $w\in\operatorname{MM}(G)$ and $t\leq\lfloor n/2\rfloor$. For all $c\in\mathcal{C}$ such that $\mu(w,c)\geq n-t$, if we have $\mu_{\mathcal{X}}(w,c)=n-t$, then for all completions $\mathcal{X}^{\prime}$ of $\mathcal{X}$, we have $\mu_{\mathcal{X}^{\prime}}(w,c)\geq n-t\geq t$. For all the other candidates $c\in\mathcal{C}$ such that $\mu(w,c)<n-t$, we take $\mu_{\mathcal{X}}(w,c)=t$ and $\mu_{\mathcal{X}}(c^{\prime},c)=n-t$ for some candidate $c^{\prime}\in\mathcal{C}$ such that $\mu(c^{\prime},c)\geq n-t$. Then for all completions $\mathcal{X}^{\prime}$ of $\mathcal{X}$, we have $\mu_{\mathcal{X}^{\prime}}(w,c)\geq t$ and $\mu_{\mathcal{X}^{\prime}}(c^{\prime},c)\geq n-t$. Let $t=\min(\sigma_{w},\left\lfloor n/2\right\rfloor)$, for all losing candidates $c\in\mathcal{C}\setminus{\{w\}}$, either $\mu(w,c)\geq n-t$ and only ntn-t pairwise comparisons are needed else n=nt+tn=n-t+t comparisons are needed. With $k=|\{c:c\in\mathcal{C},\,\mu(w,c)\geq n-t\}|$, we have $|\mathcal{X}|=(n-t)k+n(m-1-k)=n(m-1)-tk$. ∎
We provide tight bounds on the size of an SMS for $\operatorname{MM}$.
@@ 4.4 Borda and Copeland Rules
The Borda rule ($\operatorname{BO}$) chooses the set of candidates with a maximal total amount of wins across all pairwise comparisons with the other candidates.
We show that the size of an SMS for $\operatorname{BO}$ only depends on the Borda score of the winner and its worst performance against the other candidates.
Theorem 15.
@@
From the characterization of the size of SMSs for $\operatorname{BO}$, we derive tight upper and lower bounds on SMSs size.
@@
Given $n$ voters, a partial $n$-weighted tournament $G=(\mathcal{C},\mu)$ with $|\mathcal{C}|=m$, and a winning candidate $w\in\operatorname{BO}(G)$. For all SMSs $\mathcal{X}$ for $w\in\operatorname{BO}(G)$, we have $\left\lceil\frac{n(m-1)^{2}}{m}\right\rceil\leq|\mathcal{X}|\leq(m-1)\left\lfloor n\frac{m-1}{2}\right\rfloor$.
@@
Finally, we show that all SMSs share a similar structure centered around the winner’s wins by bounding the difference in terms of win count between SMSs.
@@
Since the Copeland rule ($\operatorname{CO}$) coincides with the Borda rule on 1-weighted tournaments, the following results are a special case of the previous results obtained on Borda.
@@
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$ and a winning candidate $w\in\operatorname{CO}(G)$ with a Copeland score of $\sigma_{w}$, for all SMSs $\mathcal{X}$ for $w\in\operatorname{CO}(G)$:
@@
Given a complete tournament $G=(\mathcal{C},E)$ with $|\mathcal{C}|=m$ and a winning candidate $w\in\operatorname{CO}(G)$, for all SMSs $\mathcal{X}$ for $w\in\operatorname{CO}(G)$ $m-1\leq|\mathcal{X}|\leq(m-1)\left\lfloor\frac{m-1}{2}\right\rfloor$.
@@
  • $a$ is not covered by $b$ since $a$ is preferred to $b$
  • $a$ is not covered by $c$ since $a$ is preferred to $b$ and $b$ is preferred to $c$
  • $a$ is not covered by $d$ since $a$ is preferred to $d$
@@
  • $a$ wins at least 2 pairwise comparisons in each head-to-head
  • $b$ wins at most 2=5-3 pairwise comparisons in one head-to-head
  • $c$ wins at most 2=5-3 pairwise comparisons in one head-to-head
  • $d$ wins at most 2=5-3 pairwise comparisons in one head-to-head
Figure 4: From a tournament (a) and a weighted tournament (e), we compute a smallest minimal support for $a\in\operatorname{UC}(G)$ (b) and for $a\in\operatorname{MM}(G)$ (f). We identify and visualize their underlying structure as a $a$-rooted tree (c) or by mean of the out-going edges in the neighborhood of the winning candidate and the in-going edges in those of the losing candidates (g). Finally, we produce textual explanations based on these structures (d) and (h). Examples for the remaining tournament solutions are available in Section B of the Appendix.
5 Certified User-Friendly Explanations
We now inspect MSs in light of the four Gricean principles underpinning effective explanations that we sketched in the introduction: quantity, quality, relation and manner (grice1975logic). The formal grounding of MSs ensure that the reasoning is sound, thus covering quality. By definition, MSs provide the minimal information necessary to ensure that a designated candidate is among the election winners. Thus, MSs are informative enough to support our goal and verify both the quantity and relation principles. Regarding manner, or the way information is delivered, it is possible to create explanations containing the smallest amount of information necessary thanks to our algorithms efficiently computing the smallest MSs for five out of the six studied solutions.
However, MSs still suffer from two flaws that limit their explanatory power. First, MSs are (multi-)sets of pairwise comparisons, with a flat structure that makes it hard to gain insights from the unfolding of the tournament solution. Second, their formal nature renders them more suited to be handled by machines or by domain experts than end-users. We propose to address these shortcomings by capitalizing on our structural results for MSs to organize the explanation, and by showcasing how to automatically produce intuitive textual explanation from MSs.
Figure 4 presents two examples of this process. For path-based rules such as $\operatorname{UC}$, MSs can be presented as a rooted out-tree encompassing all directed paths from the winner to the other candidates (Figure 4(c)), with each path corresponding to a bullet point in the textual explanation (see Figure 4(d)). For myopic-score rules such as $\operatorname{MM}$, MSs can present the neighborhoods of each candidate (see Figure 4(g)) from which candidates’ scores can be bounded to create a compact textual explanation (see Figure 4(h)).
@@ 6 Conclusions
Minimal supports are compact structures that can serve as certified explanations for tournament solutions. In this paper we have characterized the structure and given the size of smallest minimal supports for well-known tournament solutions, and provided polynomial algorithms for their computation for all but one case. By exploiting our characterizations we have also provided rigorous intuitive visual and textual presentations of a minimal support to demonstrate their applicability as human-oriented explanations.
Minimal supports also have far-reaching connections with several classical problems in computational social choice. In contrast with previous results on the margin of victory (brill2020refining; doring2023margin) for the top cycle, the uncovered set, Copeland, and Borda, we prove that the size of a smallest minimal support is the same for every candidate in the winner set, preventing its use as a refinement of those tournament solutions. From a query complexity perspective, while the top cycle and the uncovered set require $\Omega(m^{2})$ queries (MaitiPalash2024, Theorem 3.5 & 3.8), we show that, with a perfect oracle, exactly m1m-1 queries are needed. Although finding the smallest minimal explanation is generally hard, we provide algorithms for the top cycle, the uncovered set, Copeland, maximin and Borda with complexity $\mathcal{O}(m^{2}\log n)$, which corresponds to the size of a tournament.
Several key directions merit further exploration: empirically evaluating the usefulness of the presented explanations, studying a broader range of rules, and analyzing structural properties of minimal supports. Finally, in an approach similar to that of de2000choosing, minimal supports for different rules on the same tournament could be compared from a set-theoretical perspective.
Acknowledgments
The authors thank the reviewers of AAAI26 for their constructive comments and suggestions, which helped improve this paper.
This work is funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work is supported by ERC grant 101166894 “Advancing Digital Democratic Innovation” (ADDI).